OpenStudy (anonymous):

what is the hydroxide ion concentration of a solution that has a ph of 8.57

OpenStudy (btaylor):

pH = 14 + log[OH] 8.57 = 14 + log [OH] -5.43 = log [OH] 10^-5.43 = [OH] 3.71535229 × 10-6 = [OH]

OpenStudy (anonymous):

so 3.71535229 x 10-6 ?

OpenStudy (btaylor):

yes.

OpenStudy (anonymous):

thank you i have a nother question though

OpenStudy (anonymous):

3.71535229 × 10-6 = [OH]

OpenStudy (anonymous):

What does the equation 2h2o(l) <--> h3o+(aq) + OH-(aq) represent

OpenStudy (btaylor):

it is the self-ionization of water.

OpenStudy (anonymous):

once again thank you